OpenStudy (anonymous):

write this equation in slope-intercept form: 3+4y=16x

OpenStudy (anonymous):

slope int = y=mx+b 3+4y=16x -3 -3 4y=16x-3 /4 /4 /4 y=4x-3/4 <------- TADA